# Latency Adjustable Transformer Encoder for Language Understanding

### v8

@@ IV-E Experimental Results of GPT-2
Model Wiki Text-103 LAMBADA PTB 1BWa Speedup
GPT-2 37.5 35.1 36.46 45.54 1
DistilGPT2 43.0 - - - 1.9
ZipGPT2 72.1 - - - 3.3
This Work 43.2 59.0 52.9 51.66 2.7
49.3 64.6 61.7 54.53 3.5
a The model is evaluated on 0.5% of the 1BW test set.
Table V exhibits the classification accuracy findings of the proposed method in contrast to GPT-2 and DistilGPT2. Notably, the proposed method attains an inference latency three times quicker than GPT-2, accompanied by a marginal average accuracy reduction of 0.55%. Blank fields in the tables, indicate instances where some works did not report their assessments on certain datasets.
TABLE V: Accuracy comparison between GPT-2 [71] and DistilGPT2 [46] with latency adjustable GPT-2 (This Work)
Model SST-2 IMDB Avg Drop Speedup
GPT-2 93.2 94.1 0 1
DistilGPT2 91.5 - 1.7 1.9
This Work 92.8 93.4 0.55 3.0
IV-F Experimental Results of Flan-T5
To assess the proposed method under the instruction tuning paradigm, the method is applied to the base and large versions of the Flan-T5 model which have 220M and 770M parameters respectively. The method’s performance is evaluated using different types of prompting with the MMLU and BBH benchmarks. In the evaluation procedure, no training or fine-tuning steps are required. The modified model with the original Flan-T5 model weights is studied in this section. Table VI presents the experimental results for 5-shot and 3-shot direct prompting on the MMLU and BBH benchmarks, measuring the exact-match score. As shown in the table, the model achieves up to a 3-fold encoder speedup in the MMLU benchmark with only marginal decrease in score, and in the BBH benchmark, the method functions effectively with up to a 1.5-fold encoder speedup.
TABLE VI: Experimental results of the proposed method on Flan-T5 [72] are evaluated using the MMLU [78] and BBH [79] instruction tuning datasets.
Model MMLUa BBHb
Score Encoder Speedupc Score Encoder Speedup
Flan-T5 Base (220M) $\alpha_{ER}$ = 1.00 34.05 1 30.79 1
$\alpha_{ER}$ = 0.98 34.29 1.18 30.90 1.24
$\alpha_{ER}$ = 0.95 34.13 1.41 30.37 1.46
$\alpha_{ER}$ = 0.93 34.30 1.58 28.55 1.58
$\alpha_{ER}$ = 0.90 34.13 1.83 18.92 1.83
$\alpha_{ER}$ = 0.85 33.95 2.34 - -
$\alpha_{ER}$ = 0.80 32.24 2.94 - -
Flan-T5 Large (770M) $\alpha_{ER}$ = 1.00 41.95 1 37.04 1
$\alpha_{ER}$ = 0.98 41.92 1.29 36.48 1.45
$\alpha_{ER}$ = 0.95 42.05 1.78 29.25 1.90
$\alpha_{ER}$ = 0.93 41.51 2.23 22.81 2.37
$\alpha_{ER}$ = 0.90 35.66 3.08 16.78 3.03
@@ IV-F Experimental Results of Flan-T5
b BBH includes 23 challenging tasks, uses 3-shot direct prompting, and measure the exact-match score.
c The proposed self-attention layer is applied to the encoder stage of the Flan-T5 model, and thus the speedup of the encoder stage is reported.
@@ IV-G Offline-Tuning Property
After the fine-tuning phase, the trade-off between inference accuracy and speed of the model can be adjusted with the offline-tuning property. This property controls each encoder layer’s word-vector elimination rate by adjusting the Speedup-Coefficient hyper-parameter ($\alpha_{SC}$) value. Fig. 6 presents the offline-tuning property. In the following experiments, the model first fine-tuned with $\alpha_{SC}$ equal to one for a given task, and then in the offline-tuning phase, the $\alpha_{SC}$ value changes from 0.85 to 1.2 to observe its effect on inference speedup and accuracy.
Fig. 6a shows the model inference speedup for a range of $\alpha_{SC}$ hyper-parameter values on multiple tasks. The model inference speedup increased more than three times compared to BERTbase, with the $\alpha_{SC}$ value decreasing to 0.85. Fig. 6b presents the accuracy stability on multiple tasks. It shows that the inference results are kept stable over a wide range of $\alpha_{SC}$ values selections. Fig. 6c shows that the inference results are almost unchanged even if the complexity of the model reduces by 70% according to the proposed method. In these experiments, the model’s accuracy on validation sets is reported because of the GLUE test set verification limitations.
Figure 6: The proposed offline-tuning results, (a)Inference speedup compare to BERTbase versus $\alpha_{SC}$, (b)Validation set results versus $\alpha_{SC}$, (c)Validation set results versus speedup compare to BERTbase for MRPC [65], SST-2 [64], QNLI [67], IMDB [62] and MNLI-M/MM [69] tasks.
TABLE VII: Ablation study on the proposed self-attention layer with different self-attention Sort layer implementations.
@@
BERTbase Self-Attention 93.5 88.7 84.7 83.8 84.3 93.7 88.12
@@
a F1-score for MRPC; Pearson Correlation for STS-B; Accuracy for the rest.
The stability of the proposed method’s offline perplexity (PPL) in text generation tasks is evaluated by training the model with a specific speedup coefficient and subsequently altering this value during the inference stage, measuring the model’s zero-shot perplexity. Fig. 7 illustrates the offline perplexity of the proposed method on various datasets, including WikiText-103, LAMBADA, PTB, and 1BW. The graph depicts the model’s robust stability, with a smooth perplexity curve observed across a broad range of speedup values. This underscores the efficacy of the proposed method in maintaining stable performance under varying conditions.
IV-H Ablation Studies
Table VII presents the ablation study on the proposed self-attention layer. The first row of the table presents the original BERTbase results, and the subsequent rows are the results of the self-attention layer with different types of word-vector sorting strategies.
@@ IV-H Ablation Studies
As shown in the table, the first strategy shows a significant accuracy drop compared to the original BERTbase and the proposed method. The second strategy has a destructive effect on the model results, and the fine-tuning process becomes completely unstable in CoLA, QQP, QNLI, and RTE datasets not presented in the table. The third strategy is superior to other investigated strategies and proves that the proposed self-attention Sort layer is effective. In these experiments, the hyper-parameters values and the fine-tuning conditions are the same.
@@ V Conclusion
This paper introduces a latency-adjustable Transformer encoder layer that significantly reduces the inference time of Transformer-based architectures with only a marginal decrease in accuracy. A novel Attention Context Contribution (ACC) metric is proposed to assess context contribution in each Transformer encoder layer, experimentally demonstrating that word-vector contribution diminishes in the last encoder layers. An accurate analytical speedup estimation is presented, aiding system designers in hyper-parameter selection. The powerful offline-tuning property enables adjustment of model speedup after the fine-tuning phase. The proposed method is applied to BERTbase, GPT-2, and Flan-T5 models, validated on the GLUE benchmark, the IMDB dataset, and various text generation tasks. Additionally, the method’s ability to understand language is evaluated under the instruction tuning paradigm, employing different types of prompting to assess its versatility and robustness. The suggested approach asserts that in Large Language Models (LLMs), although the complete network is vital for training, it can be truncated during the fine-tuning phase.

### v9

@@ IV-E Experimental Results on Text Generation Tasks
Model Wiki Text-103a LAMBADA PTB 1BWb Speedup
GPT-2 37.5 35.1 36.46 45.54 1.0x
DistilGPT2 43.0 70.43 63.57 75.40 1.9x
ZipGPT2 72.1 - - - 3.3x
This Work 43.2 59.0 52.9 51.66 2.7x
49.3 64.6 61.7 54.53 3.5x
a Lower perplexity is better.
b The models are evaluated on 0.5% of the 1BW test set.
IV-F Experimental Results on Instruction Tuning Tasks
For the instruction-tuning evaluation, the Flan-T5, Gemma2, Mistral, and Llama3 models are selected. In these models, the self-attention layers are replaced with the proposed encoder self-attention layer, and the default weights are used without any additional tuning or fine-tuning. The method’s performance is evaluated using various prompting types on the MMLU benchmark. Table VI presents the experimental results for 5-shot direct prompting on the MMLU benchmark. As shown, the suggested method achieves up to a 3-fold reduction in TTFT with only a marginal decrease in score.
TABLE VI: Time-to-First-Token (TTFT) Speedup Results of the Proposed Approach on Flan-T5 [86], Gemma2 [75], Mistral [72], and Llama3 [76] Evaluated with the MMLU [92]
Instruction Tuning Dataset.
Architecture Name Model Model Size MMLU Scorea TTFT Speedupb
T5 Flan-T5-base 220M 34.05 1.00x
32.24 1.54x
Flan-T5-large 770M 41.95 1.00x
41.51 1.95x
Gemma Gemma2 2B 56.74 1.00x
56.20 1.63x
Mistral Mistral 7B 41.85 1.00x
40.83 1.65x
Llama Llama3 8B 65.69 1.00x
64.91 2.86x
@@ IV-F Experimental Results on Instruction Tuning Tasks
b The proposed attention layer is applied to the models, and the Time-to-First-Token (TTFT) speedup is reported.
@@ IV-G Offline-Tuning Property
After the fine-tuning phase, the trade-off between inference accuracy and speed of the model can be adjusted with the offline-tuning property. This property controls each encoder layer’s word-vector elimination rate by adjusting the Speedup-Coefficient hyper-parameter ($\alpha_{SC}$) value. Fig. 6 presents the offline-tuning property results on BERT-base. In the following experiments, the model first fine-tuned with $\alpha_{SC}$ equal to one for a given task, and then in the offline-tuning phase, the $\alpha_{SC}$ value changes from 0.85 to 1.2 to observe its effect on inference speedup and accuracy.
Fig. 6a shows the model inference speedup for a range of $\alpha_{SC}$ hyper-parameter values on multiple tasks. The model inference speedup increased more than three times compared to BERT-base, with the $\alpha_{SC}$ value decreasing to 0.85. Fig. 6b presents the accuracy stability on multiple tasks. It shows that the inference results are kept stable over a wide range of $\alpha_{SC}$ values selections. Fig. 6c shows that the inference results are almost unchanged even if the complexity of the model reduces by 70% according to the proposed method. In these experiments, the model’s accuracy on validation sets is reported because of the GLUE test set verification limitations.
Figure 6: The proposed offline-tuning results, (a)Inference speedup compare to BERT-base versus $\alpha_{SC}$, (b)Validation set results versus $\alpha_{SC}$, (c)Validation set results versus speedup compare to BERT-base for MRPC [80], SST-2 [79], QNLI [82], IMDB [77] and MNLI-M/MM [84] tasks.
The offline perplexity (PPL) stability of the proposed method for text generation tasks is evaluated by training the model with a specific speedup coefficient and then modifying this value during inference, while measuring the model’s zero-shot perplexity. Fig. 7 illustrates the method’s offline perplexity results on GPT-2 across various datasets, including WikiText-103, LAMBADA, PTB, and 1BW. The graph demonstrates the model’s robust stability, showing a smooth perplexity curve across a wide range of speedup values, which highlights the solution’s effectiveness in maintaining consistent performance under varying conditions.
IV-H Ablation Studies
Table VII presents the ablation study on the proposed self-attention layer. The first row of the table presents the original BERT-base results, and the subsequent rows are the results of the self-attention layer with different types of word-vector sorting strategies.
TABLE VII: Ablation study on the proposed self-attention layer with different self-attention Sort layer
implementations strategies on BERT-base [3].
@@ IV-H Ablation Studies
BERT-base Self-Attention 93.5 88.7 84.7 83.8 84.3 93.7 88.12
@@
a F1-score for MRPC, Pearson Correlation for STS-B, and Accuracy for the remaining tasks.
@@
As shown in the table, the first strategy shows a significant accuracy drop compared to the original BERT-base and the proposed method. The second strategy has a destructive effect on the model results, and the fine-tuning process becomes completely unstable in CoLA, QQP, QNLI, and RTE datasets not presented in the table. The third strategy is superior to other investigated strategies and proves that the proposed self-attention Sort layer is effective. In these experiments, the hyper-parameters values and the fine-tuning conditions are the same.
@@ V Conclusion
This paper presents a latency-adjustable Transformer architecture applicable to a wide range of models, with a marginal decrease in accuracy. We introduce a novel Attention Context Contribution (ACC) metric to evaluate the context contribution of each Transformer encoder layer, demonstrating experimentally that word-vector contributions diminish in the last encoder layers. Extensive experiments on language understanding, text generation, and instruction-tuning tasks validate the suggested technique’s effectiveness across diverse datasets, with minimal impact on the input’s global context. The paper also provides an accurate analytical estimation of speedup, assisting system designers with hyper-parameter selection. Additionally, the method’s powerful offline-tuning capability allows for model speedup adjustments after the fine-tuning phase. The proposed approach highlights that, while the complete network is essential for training, it can be truncated during the fine-tuning phase in Large Language Models (LLMs).